# PixelVLA: Advancing Pixel-level Understanding in Vision-Language-Action Model

### v1

@@ Appendix A Appendix
Figure 5: The expisode example in our Pixel-160K dataset.
Figure 6: The proposed automated annotation pipeline containing for generating visual prompts and mask annotations at scale for a given robot dataset, consisting of a gripper-aware region proposal stage and a multimodal object segmentation stage.
@@ A.1 Pixel-160K dataset
6. Annotations Generation: For the $\eta$-th episode, we feed the target object text together with the region proposal $\mathbf{R}_{\eta}$} into an open-vocabulary detection and segmentation pipeline comprising Grounding DINO (Liu et al. (2024)) and SAM (Kirillov et al. (2023)). Grounding DINO detects candidate object bounding boxes conditioned on the target object text, grounding the language query into specific image regions. Subsequently, SAM refines these detections by generating pixel-level segmentation masks for the grounded boxes. The outputs include bounding boxes, segmentation masks, and text-object alignments. We then filter the results by selecting the mask with the highest confidence score within the region proposal, discarding low-confidence or irrelevant detections. Finally, from the retained object mask, we derive visual prompts by: randomly sampling points inside the mask, generating random lines within the object area, and extracting external bounding boxes via contour detection.
@@ A.2 Experiment setup
Evaluation tasks. As shown in Fig. 7, we conduct all experiments on three simulation benchmarks, i.e., SimplerEnv-Google Robot (Li et al. (2024b)), SimplerEnv-WidowX (Li et al. (2024b)) and LIBERO (Liu et al. (2023a)). Specifically, SimplerEnv Li et al. (2024b) is an open-source simulation suite designed for reproducible and scalable evaluation of robot manipulation policies. It explicitly addresses the visual and dynamic gaps between simulation and real hardware, enabling more faithful assessments of policy generalization. SimplerEnv provides two complementary evaluation settings: Visual Matching (VM), which minimizes visual appearance discrepancies to improve the correlation between simulation and real-world performance; and Variant Aggregation (VA), which, inspired by domain randomization, introduces diverse visual perturbations and aggregates results across multiple randomized environments to obtain more robust performance estimates. Building on these capabilities, we perform zero-shot object manipulation experiments on SimplerEnv to benchmark the effectiveness of our approach.
@@
Figure 7: Experimental Setup Overview. We evaluate PixelVLA on three simulation benchmarks: SimpleEnv-Bridge with WidowX, SIMPLER-Google Robot with Google Robot, and LIBERO with Franka.
@@ A.3 Implementation details
Training. In the first stage of continuous action training, we initialize PixelVLA’s vision encoder, MLP projector, and LLM backbone with pretrained weights from VLAs (Kim et al. (2024); Black et al. (2024)), which were trained on the large-scale OXE dataset (O’Neill et al. (2024)). To focus on learning continuous action mappings, the visual prompting encoder and the multiscale pixel-aware encoder are removed, while all other modules except the continuous action decoder are frozen to preserve general manipulation knowledge. The continuous action decoder maps the final hidden states of the LLM to continuous action values. We use L1 regression (Zhao et al. (2023); Kim et al. (2025)) to align the predicted actions with ground-truth actions. Each continuous action dimension is uniformly discretized into 256 bins and normalized to [1,+1][-1,+1] for fine-grained prediction. The input observation consists of a single third-person camera view resized to $224\times 224$ pixels. Actions are predicted in chunks of 8 timesteps, with $\mathbf{a^{t}}\in\mathbb{R}^{8\times 7}$. During this stage, PixelVLA is trained on a mixture of Fractal (Brohan et al. (2022)) and Bridge v2 (Walke et al. (2023)) datasets. Training runs for 100k steps with a batch size of 32 and a learning rate of $5\times 10^{-4}$. This stage is omitted when adapting PixelVLA to $\pi_{0}$ due to the effectiveness of its pretrained action expert.
In Pixel-level Understanding Enhancement Stage, PixelVLA is fine-tuned for pixel-level understanding using the Pixel-160K dataset. The visual prompting encoder and multiscale pixel-aware encoder are jointly trained, while LoRA adaptation is applied to the LLM backbone (rank $r=32$). All other modules remain frozen except the continuous action decoder, which is optimized using L1 regression, following the same continuous action representation as in the first stage. Input observations are a single third-person camera view resized to $224\times 224$ pixels. Actions are predicted in chunks of 8 timesteps, $\mathbf{a^{t}}\in\mathbb{R}^{8\times 7}$. Training runs for 200k steps with a batch size of 32 and a learning rate of $1\times 10^{-3}$, enhancing PixelVLA’s ability to localize and manipulate target objects at the pixel level.
@@ A.4 Qualitative comparisons
Table 5: SimplerEnv (Li et al. (2024b)) simulation valuation results in terms of the average success rate for the Google Robot setup. VM denotes Visual Matching and VA is Variant Aggregation. ◼ denotes tuning-based methods applied to the pretrained weights of OpenVLA.

### v2

@@ Appendix A Appendix
(5) Limitations in Sec. A.5.
Figure 5: The episode example in our Pixel-160K dataset.
Figure 6: The proposed automated annotation pipeline for generating visual prompts and mask annotations at scale for a given robot dataset, consisting of a gripper-aware region proposal stage and a multimodal object segmentation stage.
@@ A.1 Pixel-160K dataset
6. Annotations Generation: For the $\eta$-th episode, we feed the target object text together with the region proposal $\mathbf{R}_{\eta}$} into an open-vocabulary detection and segmentation pipeline comprising Grounding DINO (Liu et al. (2024a)) and SAM (Kirillov et al. (2023)). Grounding DINO detects candidate object bounding boxes conditioned on the target object text, grounding the language query into specific image regions. Subsequently, SAM refines these detections by generating pixel-level segmentation masks for the grounded boxes. The outputs include bounding boxes, segmentation masks, and text-object alignments. We then filter the results by selecting the mask with the highest confidence score within the region proposal, discarding low-confidence or irrelevant detections. Finally, from the retained object mask, we derive visual prompts by: randomly sampling points inside the mask, generating random lines within the object area, and extracting external bounding boxes via contour detection.
@@ A.2 Experiment setup
Evaluation tasks. As shown in Fig. 7, we conduct all experiments on three simulation benchmarks, i.e., SimplerEnv-Google Robot (Li et al. (2024c)), SimplerEnv-WidowX (Li et al. (2024c)) and LIBERO (Liu et al. (2023a)). Specifically, SimplerEnv Li et al. (2024c) is an open-source simulation suite designed for reproducible and scalable evaluation of robot manipulation policies. It explicitly addresses the visual and dynamic gaps between simulation and real hardware, enabling more faithful assessments of policy generalization. SimplerEnv provides two complementary evaluation settings: Visual Matching (VM), which minimizes visual appearance discrepancies to improve the correlation between simulation and real-world performance; and Variant Aggregation (VA), which, inspired by domain randomization, introduces diverse visual perturbations and aggregates results across multiple randomized environments to obtain more robust performance estimates. Building on these capabilities, we perform zero-shot object manipulation experiments on SimplerEnv to benchmark the effectiveness of our approach.
@@
Figure 7: Experimental Setup Overview. We evaluate PixelVLA on three simulation benchmarks: SimpleEnv-WidowX with WidowX robot, SimpleEnv-Google Robot with Google robot, and LIBERO with Franka.
@@ A.3 Implementation details
Training. In the first stage of continuous action training, we initialize PixelVLA’s vision encoder, MLP projector, and LLM backbone with pretrained weights from VLAs (Kim et al. (2024); Black et al. (2024)), which were trained on the large-scale OXE dataset (O’Neill et al. (2024)). To focus on learning continuous action mappings, the visual prompt-aware encoder and the multiscale pixel-aware encoder are removed, while all other modules except the continuous action decoder are frozen to preserve general manipulation knowledge. The continuous action decoder maps the final hidden states of the LLM to continuous action values. We use L1 regression (Zhao et al. (2023); Kim et al. (2025)) to align the predicted actions with ground-truth actions. Each continuous action dimension is uniformly discretized into 256 bins and normalized to [1,+1][-1,+1] for fine-grained prediction. The input observation consists of a single third-person camera view resized to $224\times 224$ pixels. Actions are predicted in chunks of 8 timesteps, with $\mathbf{a^{t}}\in\mathbb{R}^{8\times 7}$. During this stage, PixelVLA is trained on a mixture of Fractal (Brohan et al. (2022)) and Bridge v2 (Walke et al. (2023)) datasets. Training runs for 100k steps with a batch size of 32 and a learning rate of $5\times 10^{-4}$. This stage is omitted when adapting PixelVLA to $\pi_{0}$ due to the effectiveness of its pretrained action expert.
In Pixel-level Understanding Enhancement Stage, PixelVLA is fine-tuned for pixel-level understanding using the Pixel-160K dataset. The visual prompt-aware encoder and multiscale pixel-aware encoder are jointly trained, while LoRA adaptation is applied to the LLM backbone (rank $r=32$). All other modules remain frozen except the continuous action decoder, which is optimized using L1 regression, following the same continuous action representation as in the first stage. Input observations are a single third-person camera view resized to $224\times 224$ pixels. Actions are predicted in chunks of 8 timesteps, $\mathbf{a^{t}}\in\mathbb{R}^{8\times 7}$. Training runs for 200k steps with a batch size of 32 and a learning rate of $1\times 10^{-3}$, enhancing PixelVLA’s ability to localize and manipulate target objects at the pixel level.
@@ A.4 Qualitative comparisons
Table 5: SimplerEnv (Li et al. (2024c)) simulation evaluation results in terms of the average success rate for the Google Robot setup. VM denotes Visual Matching and VA is Variant Aggregation. ◼ denotes tuning-based methods applied to the pretrained weights of OpenVLA.
@@
A.5 Limitations
Although PixelVLA substantially improves VLA performance through pixel-level understanding and tailored visual prompts, it remains limited in handling richer input modalities (e.g., 3D perception) and more advanced forms of visual prompting, such as precise trajectory guidance, reference-image prompts, pose-conditioned prompts, or compositional prompt sequences. In addition, while PixelVLA is extensively validated across three simulated benchmarks, we expect that real-world robot experiments would further strengthen its contributions. We regard these extensions as important directions for future work.